OpenStudy (anonymous):

Is x = 3 the only solution to this:

OpenStudy (anonymous):

to what??

OpenStudy (anonymous):

\[\sqrt{2x+3} - \sqrt{x + 1} = 1\]

OpenStudy (anonymous):

to this

OpenStudy (anonymous):

it's 3 and -1 -1 is also there

OpenStudy (anonymous):

at last u will get this equation x^2-2x-3=0 solve for x and get x=3 and x=-1